# ShowHowTo: Generating Scene-Conditioned Step-by-Step Visual Instructions

### v1

@@ Appendix A Dataset Collection Details
Step extraction details. Figure 20 illustrates our Llama 3 prompt for extracting instructional steps from video narrations, including the few-shot examples used (truncated to fit into a page; complete prompts are available on the project website). We processed videos under 10 minutes long, as longer transcripts exceeded the context limit of Llama 3 and also degraded output quality. The model is instructed to generate temporally ordered steps with approximate timestamps using video transcripts. We filter out malformed results with non-temporally ordered steps or incomplete descriptions. Examples of extracted steps are shown in Figure 8.
@@
To account for any misalignment of narrations with respect to the video frames, we expand the temporal boundaries by a fixed duration of $\epsilon$ seconds, increasing the search space for frame alignment. Through analysis on a small validation set of manually annotated videos with precise step boundaries, we found $\epsilon=15$ seconds yielded a good balance of precision (the correct frame was selected by DFN-CLIP) and recall (the correct frame was inside the searched interval).
We use DFN-CLIP over related contrastive models such as CLIP [47] and SigLIP [66] as we found it superior in certain cases for matching video frames to instructional steps. We observed CLIP and SigLIP often exhibited limitations such as incorrect object state identification (e.g., unroasted peppers, uncut prosciutto in Figure 7, rows 1-2) and tendency to select blurry or transition frames (Figure 7, rows 2-3).
@@ Appendix E Additional Quantitative Results
Downstream applications of ShowHowTo. In GenHowTo [53], the authors propose to evaluate generative methods in a downstream application, where the method generates images of various classes that are then used to train a simple classifier. The performance of this classifier is then computed on the real set of images. We show the results of our method in Table 7. We evaluate our model as is, i.e., trained on the ShowHowTo dataset without any additional training or fine-tuning. We report the action accuracy metric $Acc_{ac}$, which evaluates whether our generated visual instruction images can be used for downstream application of classifying actions. Note that this metric is image-based and does not evaluate sequences. Nonetheless, ShowHowTo improves over the previous state-of-the-art from [53] by 6 percentage points.
@@
Training data source Accac
@@
Table 7: Performance of classifiers trained on generated images according to the GenHowTo protocol [53]. ShowHowTo model outperforms the prior state-of-the-art with no fine-tuning on the GenHowTo train set.

### v2

@@ Appendix A Dataset Collection Details
Step extraction details. Our approach uses Llama 3 to extract the steps from video narrations. This contrasts the related work [53], which extracts steps from video frames and uses an image-captioning model to generate step captions. We found that the latter approach resulted in captions that were too brief, high-level, and lacked sufficient detail to differentiate adjacent steps.
Figure 20 illustrates our Llama 3 prompt for extracting instructional steps from video narrations, including the few-shot examples used (truncated to fit into a page; complete prompts are available on the project website). We processed videos under 10 minutes long, as longer transcripts exceeded the context limit of Llama 3 and also degraded output quality. The model is instructed to generate temporally ordered steps with approximate timestamps using video transcripts. We filter out malformed results with non-temporally ordered steps or incomplete descriptions. Examples of extracted steps are shown in Figure 8.
@@
Although the temporal boundaries generated by Llama 3 are rarely incorrect, narrations do not always align with respect to video frames and can occur slightly before/after the visuals. To account for the well-known temporal misalignment [25, 37], we expand the temporal boundaries by a fixed duration of $\epsilon$ seconds, increasing the search space for frame alignment. Through analysis on a small validation set of manually annotated videos with precise step boundaries, we found $\epsilon=15$ seconds yielded a good balance of precision (the correct frame was selected by DFN-CLIP) and recall (the correct frame was inside the searched interval).
We use DFN-CLIP over related contrastive models such as CLIP [47] and SigLIP [66] as we found it superior in certain cases for matching video frames to instructional steps. We observed CLIP and SigLIP often exhibited limitations such as incorrect object state identification (e.g., unroasted peppers, uncut prosciutto in Figure 7, rows 1-2) and tendency to select blurry or transition frames (Figure 7, rows 2-3). We quantify the quality of the automatically selected keyframes by a small user study. We manually annotated keyframes for 100 steps and asked humans to blindly select whether the manually or automatically selected frame better corresponds to the text instruction. DFN-CLIP was preferred in 18% of cases, human annotation was preferred in 36% of cases, remaining 46% of cases were a tie. This indicates fairly decent alignment with human judgment.
@@ Appendix E Additional Quantitative Results
Method Accac
@@
Table 7: Zero-shot evaluation on the GenHowTo dataset according to the GenHowTo protocol [53]. The ShowHowTo model outperforms the prior state-of-the-art with no fine-tuning on the GenHowTo train set.
Zero-shot evaluation on the GenHowTo dataset. In GenHowTo [53], the authors propose to evaluate generative methods in a downstream application, where the method generates images of various classes that are then used to train a simple classifier. The performance of this classifier is then computed on the real set of images. We show the results of our method in Table 7. We evaluate our model as is, i.e., trained on the ShowHowTo dataset without any additional training or fine-tuning. We report the action accuracy metric $Acc_{ac}$, which evaluates whether our generated visual instruction images can be used for downstream application of classifying actions. Note that this metric is image-based and does not evaluate sequences. Nonetheless, ShowHowTo improves over the previous state-of-the-art from [53] by 6 percentage points.
Number of generated frames
1 2 3 4 5 6 7
Step Faith. 1.00 0.72 0.59 0.52 0.50 0.50 0.51
Scene Consist. 0.20 0.29 0.43 0.42 0.34 0.36 0.31
Task Faith. 0.30 0.49 0.48 0.45 0.42 0.46 0.37
Table 8: Model’s performance for different lengths of the generated sequence. The performance is fairly similar across different lengths, with a decrease observed for longer sequences.
Variable sequence length generation analysis. Our model generates sequences of variable length. We analyze the performance of the model for various sequence lengths in Table 8. Except for the degenerative case with one frame, the performance is fairly similar across different lengths, with a decrease observed for longer sequences.
Task category Step Scene Task Average
Faithf. Consist. Faithf.
Cars & Other Vehicles 0.34 0.37 0.58 0.43
Education and Communications 0.37 0.55 0.24 0.39
Food and Entertaining 0.67 0.24 0.40 0.44
Health 0.29 0.56 0.46 0.44
Hobbies and Crafts 0.39 0.46 0.41 0.42
Holidays and Traditions 0.56 0.35 0.52 0.48
Home and Garden 0.40 0.41 0.38 0.40
Pets and Animals 0.48 0.41 0.39 0.43
Sports and Fitness 0.36 0.42 0.55 0.44
Table 9: Model’s performance for different task categories. Task categories with only a few sequences not shown.
Per-task performance analysis. We report the performance of our model in various HowTo100M task categories in Table 9. We observe that the performance is significantly dependent on the task distribution in the dataset. In detail, for the Step Faithfulness metric, the best performance is achieved in cooking tasks because of plentiful training data and clear, visually distinct steps. On the other hand, for Scene Consistency, the cooking tasks perform the worst as the matching is done across the whole dataset, where a large portion is cooking with many similar scenes and frames. Additionally, the cooking tasks contain many close-up scenes without any scene background that can be used for matching the generated images to the correct sequence. The best Scene Consistency is achieved for tasks in the Health and the Education categories due to the uniqueness of the sequences in those categories.